Identify the focus and directrix of the following parabola
(a) y^2 = 3x^2
1
Expert's answer
2021-01-27T06:31:27-0500

Given:

Identify the focus and directrix of the following parabola

(a) y^2 = 3x^2


Solution:


1) It is not a parabola


Possible solutions:


2) If there is a mistake in the left part of the equation and it should be like:

"y = 3x^2"

Then, remembering the normal form of parabola:

"(x-h)^2 = 4p(y-k)"

Reshaping our equation:

"(x - 0)^2 = 4 * 1\/4 * 1\/3 * (y-0)"

"(x-0)^2=4*1\/12(y-0)"

From which:

"h=0;p=1\/12;k=0"

Focus and directrix are:

"focus=(h, k+p); directrix:y=k-p"

In our case:

"focus=(0, 1\/12); directrix: y=-1\/12"


3) If there is a mistake in the right part of the equation and it should be like:

"y^2=3x"

The normal form of parabola in that case:

"(y-k)^2=4p(x-h)"

In our case:

"(y-0)^2=4 * 1\/4*3x^2"

"(y-0)^2=4*3\/4*x^2"

From which:

"k=0;p=3\/4;h=0"

Focus and directrix are:

"focus=(h+p,k); directrix: x=h-p"

In our case:

"focus = (3\/4,0); directrix: x=-3\/4"


Answer:

It is not a parabola